A carpenter has three pieces of wood measuring 210cm,245cm, and 315'cm in lenght. She wants to cut them into shorter pieces that are all the same length and with no wood left over. what is the greatest possible lenght of each shorted piece?
Answer by stanbon(75887)

A carpenter has three pieces of wood measuring 210cm,245cm, and 315'cm in lenght. She wants to cut them into shorter pieces that are all the same length and with no wood left over. what is the greatest possible length of each shorted piece?
----
210 = 2*105 = 2*5*21 = 2*3*5*7
-----------------
245 = 5*49 = 5*7^2
-------------------
315 = 5*63 = 3^2 * 5* 7
-----
Greatest common Divisor:: 5*7 = 35
-----
210 = 35*6:: 6 pieces of size 35
245 = 35*7:: 7 pieces of size 35
315 = 35*9:: 9 pieces of size 35
----------
Cheers,
Stan H.